# ShortGPT: Layers in Large Language Models are More Redundant Than You Expect

### 06 Mar 2024 v1

@@ 1 Introduction
The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains. Guided by the scaling laws identified in prior work Kaplan et al. (2020); Hoffmann et al. (2022), current LLM research has tended to increase model parameters consistently to boost performance. However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use. To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques Zhu et al. (2023) that reduce the inference cost. Model compression methods can be broadly categorized into two categories: quantization Liu et al. (2021); Gholami et al. (2022); Dettmers et al. (2022, 2024) and pruning LeCun et al. (1989); Han et al. (2015). Quantization methods quantize the weights and activations to lower precision. However, the acceleration benefits of quantization are dependent on hardware support and sometimes require additional fine-tuning to maintain good performance. In contrast, pruning methods remove redundant model parameters to reduce the overall parameter count. Pruning can be directly applied to a trained model without retraining and is generally more hardware-friendly than quantization approaches. While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex. Some require the use of gradient information Zhang et al. (2023) or only focus on compression in width Ashkboos et al. (2024), rendering them overly sophisticated for practical applications. Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models. Identifying effective model pruning techniques necessitates studying model redundancy Huang et al. (2021); Dalvi et al. (2020). Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers. For model pruning in LLMs, most previous works have concentrated on tensor-wise redundancy analysis, investigating the redundancy within each parameter tensor. However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance. For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark Hendrycks et al. (2020) only drop from 55.0 to 52.22 (a 95% retention). Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.19 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model. In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs. We find BI is a more relevant indicator of the layer’s importance in LLMs, and we can perform model pruning by simply deleting redundant layers with BI. This straightforward act of removing specific layers significantly outperforms previous more complex pruning methods. Our findings highlight substantial redundancy in current LLM architectures and shed light on opportunities for more efficient LLM training in the future. The main contributions of our paper are summarized as follows:
@@
Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. 2.2), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
@@ 2.1 Layer redundancy
The predominant LLMs in current are primarily based on the Transformer Vaswani et al. (2017). The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other. Transformer is a sequence to sequence mapping, which can be defined as $y=f(X,\theta)$, where $X\in R^{T\times n}$, $Y\in R^{T\times n}$, $T$ is the length of the sequence, $n$ is the vocabulary size, $\theta$ is the learnable parameters. The formal expression of an L-layer transformer is as follows:
@@
where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ Chowdhery et al. (2023), $d$ is the hidden dim of the transformer. ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers. Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are. Previous works found that the Transformer possesses certain semantic capabilities in earlier layers Hasan et al. (2021). In this work, we uncover a significant level of redundancy between the layers of the transformer.
@@
We observed that in many cases, the network’s prediction of $Y$ remains largely unchanged. This phenomenon demonstrates a significant level of redundancy across the layers of the transformer. Figure 3 shows the perplexity and MMLU Hendrycks et al. (2020) score of omitting a certain layer in the Llama2-7B-Base Touvron et al. (2023), an English based LLMs, and Baichuan2-7B-Base Yang et al. (2023) which is mainly focused on Chinese. For more benchmark results, please refer to the section 3. Figure 3 reveals that some layers do not play a crucial role in LLMs. Dropping a certain layer from the LLMs may have a minimal effect on the final results, and the same holds for the perplexity. Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the final layer often being more critical.
@@ 2.2 Layer importance
As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers. Notably, the deeper layers demonstrate a higher level of redundancy. To remove these redundant layers, we need a model-intrinsic metric to measure the importance of the layer. An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons. In Samragh et al. (2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers. To characterize the influence of a certain layer. We defined a new metric, Block Influence (BI), under the assumption that the more a transformer block changes the hidden states, the more influential this layer is. As Figure 4 shows, the BI score of $i^{th}$ blocks can be calculated as follows:
@@ 2.3 Layer Removal
The method we propose is straightforward: layer removal, which we simply referred as ShortGPT. We delete certain layers in LLMs based on BI score. Figure 2 illustrates our pruning method. First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 Rae et al. (2019). Then we collect the hidden states of each layer during inference on these samples. Next, we calculate the BI score based on the collected hidden states. Finally, we sort layers in ascending order according to the BI, and delete the layers with the smaller importance. The number of layers to be deleted can vary to trade off the speed and performance. The details of our layer remove setting can be found in Appendix B
@@ 3.1.1 Models
To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B Touvron et al. (2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B. They are all large language models based on the decoder-only Transformer architecture. LLaMA 2 was trained on more than 2 trillion tokens. Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE Su et al. (2024) positional embedding with ALiBi Press et al. (2021).
@@ 3.1.2 Benchmarks
In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU Hendrycks et al. (2020), CMMLU Li et al. (2024) for evaluating large models. In addition, we also followed LaCo Yang et al. (2024) to evaluate a wider dataset.
MMLU Hendrycks et al. (2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings. This makes benchmarks more challenging and similar to the way we evaluate humans. This benchmark covers 57 subjects including STEM, humanities, social sciences, etc. Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
CMMLU Li et al. (2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture. CMMLU covers 67 topics, from elementary school to university or professional level. Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
CMNLI Xu et al. (2020) is part of the Chinese language understanding assessment benchmark. It consists of two parts: XNLI and MNLI. HellaSwag (HeSw) Zellers et al. (2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans. PIQA Bisk et al. (2020) is a multi-choice question and answer dataset that focuses on daily scenarios. This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios. CHID Zheng et al. (2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms. CoQA Reddy et al. (2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers. BoolQ Clark et al. (2019) is a question-answer dataset containing 15942 examples of yes/no questions. These problems occur naturally - they are generated in an environment that is silent and unconstrained. Race Lai et al. (2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students. XSumHasan et al. (2021) is used to evaluate abstract single document summarization systems. The goal is to create a short, one-sentence new summary of what the article is about. C3 Sun et al. (2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams. PG19 Rae et al. (2019) is a long document dataset from books used to test the effectiveness of language modeling.
@@ 3.1.3 Baselines
To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo Yang et al. (2024). For our method, we use PG19 for layer importance and perplexity calculation.
LLMPru. Ma et al. (2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality.
SliceGPT Ashkboos et al. (2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network.
LaCo Yang et al. (2024) is a pruning method for large language models based on merging multiple layers.
@@ 4.1 Varying prune ratio
4.2 Depth Redundancy vs width Redundancy
@@ 4.3 Redundancy on non-transformer LLM
We further investigate whether the redundancy observed in depth is a result of the Transformer’s architecture. Given that the vast majority of popular LLMs are based on the Transformer structure. We selected the RWKV-7B model proposed in Peng et al. (2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent. We analyzed this model using the methodology from 2.1. Figure 7 shows the redundancy of the RWKV-7B 111 We use rwkv-v5-world-7B from https://huggingface.co/BlinkDL/rwkv-5-world model. Through this figure, we can observe that the RWKV-7B model also exhibits a high level of redundancy. This may suggest that the redundancy is universal across LLMs.
@@ 4.4 Importance metric
Figure 8: Different importance metrics.
@@
Reverse-order/Norm: This metric posits that importance is inversely proportional to the sequence order. It assigns higher importance scores to the shallower layers Inspired by Figure 3. This method gives the same order as measuring importance by hidden states norm mentioned in 2.2. Here, the importance metric is the layer’s index itself.
Relative Magnitude: Proposed in Samragh et al. (2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
@@
Figure 8 demonstrates the different metrics. We observe that shallower layers in the LLM network are more crucial than deeper ones. Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics. The method of Samragh et al. (2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent. It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation. This also indicates the limitations of the current LLMs evaluation methods.
@@ 5 Limitation
Although our method demonstrates strong competitiveness compared to current pruning methods, with some benchmarks even suggesting that deleting certain layers has no impact on the final result, this does not imply that layer removal comes without drawbacks. Due to the limitations of benchmarking, current evaluations may not fully capture the impact of layer removal on model performance. Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks. On benchmarks such as GSM8K Cobbe et al. (2021) and HumanEval Chen et al. (2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero. Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal. Post-training techniques could potentially further mitigate performance losses, which is an area of ongoing research.

### 07 Mar 2024 v2

@@ 1 Introduction
The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains. Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research has tended to increase model parameters consistently to boost performance. However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use. To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques (Zhu et al., 2023) that reduce the inference cost. Model compression methods can be broadly categorized into two categories: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) and pruning (LeCun et al., 1989; Han et al., 2015). Quantization methods quantize the weights and activations to lower precision. However, the acceleration benefits of quantization are dependent on hardware support and sometimes require additional fine-tuning to maintain good performance. In contrast, pruning methods remove redundant model parameters to reduce the overall parameter count. Pruning can be directly applied to a trained model without retraining and is generally more hardware-friendly than quantization approaches. While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex. Some require the use of gradient information (Zhang et al., 2023) or only focus on compression in width (Ashkboos et al., 2024), rendering them overly sophisticated for practical applications. Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models. Identifying effective model pruning techniques necessitates studying model redundancy (Huang et al., 2021; Dalvi et al., 2020). Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers. For model pruning in LLMs, most previous works have concentrated on tensor-wise redundancy analysis, investigating the redundancy within each parameter tensor. However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance. For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark (Hendrycks et al., 2020) only drop from 55.0 to 52.2 (a 95% retention). Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.2 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model. In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs. We find BI is a more relevant indicator of the layer’s importance in LLMs, and we can perform model pruning by simply deleting redundant layers with BI. This straightforward act of removing specific layers significantly outperforms previous more complex pruning methods. Our findings highlight substantial redundancy in current LLM architectures and shed light on opportunities for more efficient LLM training in the future. The main contributions of our paper are summarized as follows:
@@
Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. LABEL:layerimportacne), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
@@ 2.1 Layer redundancy
The predominant LLMs in current are primarily based on the Transformer (Vaswani et al., 2017). The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other. Transformer is a sequence to sequence mapping, which can be defined as $y=f(X,\theta)$, where $X\in R^{T\times n}$, $Y\in R^{T\times n}$, $T$ is the length of the sequence, $n$ is the vocabulary size, $\theta$ is the learnable parameters. The formal expression of an L-layer transformer is as follows:
@@
where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ (Chowdhery et al., 2023), $d$ is the hidden dim of the transformer. ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers. Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are. Previous works found that the Transformer possesses certain semantic capabilities in earlier layers (Hasan et al., 2021). In this work, we uncover a significant level of redundancy between the layers of the transformer.
@@
We observed that in many cases, the network’s prediction of $Y$ remains largely unchanged. This phenomenon demonstrates a significant level of redundancy across the layers of the transformer. Figure 3 shows the perplexity and MMLU (Hendrycks et al., 2020) score of omitting a certain layer in the Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese. For more benchmark results, please refer to the section 3. Figure 3 reveals that some layers do not play a crucial role in LLMs. Dropping a certain layer from the LLMs may have a minimal effect on the final results, and the same holds for the perplexity. Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the final layer often being more critical.
@@ 2.2 Layer importance
As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers. Notably, the deeper layers demonstrate a higher level of redundancy. To remove these redundant layers, we need a model-intrinsic metric to measure the importance of the layer. An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons. In (Samragh et al., 2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers. To characterize the influence of a certain layer. We defined a new metric, Block Influence (BI), under the assumption that the more a transformer block changes the hidden states, the more influential this layer is. As Figure 4 shows, the BI score of $i^{th}$ blocks can be calculated as follows:
@@ 2.3 Layer Removal
The method we propose is straightforward: layer removal, which we simply referred as ShortGPT. We delete certain layers in LLMs based on BI score. Figure 2 illustrates our pruning method. First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 (Rae et al., 2019). Then we collect the hidden states of each layer during inference on these samples. Next, we calculate the BI score based on the collected hidden states. Finally, we sort layers in ascending order according to the BI, and delete the layers with the smaller importance. The number of layers to be deleted can vary to trade off the speed and performance. The details of our layer remove setting can be found in Appendix B
@@ 3.1.1 Models
To validate the effectiveness of our method, we conducted experiments on existing popular open-source language models, including Llama2-7B (Touvron et al., 2023), Llama2-13B, Baichuan2-7B, and Baichuan2-13B. They are all large language models based on the decoder-only Transformer architecture. LLaMA 2 was trained on more than 2 trillion tokens. Baichuan-series was mainly trained in Chinese and its 13-Billion model replaced the RoPE (Su et al., 2024) positional embedding with ALiBi (Press et al., 2021).
@@ 3.1.2 Benchmarks
In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024) for evaluating large models. In addition, we also followed LaCo (Yang et al., 2024) to evaluate a wider dataset.
MMLU (Hendrycks et al., 2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings. This makes benchmarks more challenging and similar to the way we evaluate humans. This benchmark covers 57 subjects including STEM, humanities, social sciences, etc. Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
CMMLU (Li et al., 2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture. CMMLU covers 67 topics, from elementary school to university or professional level. Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
CMNLI (Xu et al., 2020) is part of the Chinese language understanding assessment benchmark. It consists of two parts: XNLI and MNLI. HellaSwag (HeSw) (Zellers et al., 2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans. PIQA (Bisk et al., 2020) is a multi-choice question and answer dataset that focuses on daily scenarios. This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios. CHID (Zheng et al., 2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms. CoQA (Reddy et al., 2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers. BoolQ (Clark et al., 2019) is a question-answer dataset containing 15942 examples of yes/no questions. These problems occur naturally - they are generated in an environment that is silent and unconstrained. Race (Lai et al., 2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students. XSum(Hasan et al., 2021) is used to evaluate abstract single document summarization systems. The goal is to create a short, one-sentence new summary of what the article is about. C3 (Sun et al., 2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams. PG19 (Rae et al., 2019) is a long document dataset from books used to test the effectiveness of language modeling.
@@ 3.1.3 Baselines
To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo (Yang et al., 2024). For our method, we use PG19 for layer importance and perplexity calculation.
LLMPru. (Ma et al., 2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. LLMPru. applies post training to the pruned model, but for fair comparison, we do not apply post training to it.
SliceGPT (Ashkboos et al., 2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network. Specifically, they applied PCA to the hidden representation from shallow to deep layers, and incorporated the dimension reduction matrix into existing network parameters.
LaCo (Yang et al., 2024) is a pruning method for large language models based on reducing layers. LaCo gradually merges similar layers from deep to shallow and sets a threshold to avoid continuously merging too many layers.
@@ 4.1 Varying prune ratio
4.2 Depth Redundancy v.s. Width Redundancy
@@ 4.3 Redundancy on non-transformer LLM
We further investigate whether the redundancy observed in depth is a result of the Transformer’s architecture. Given that the vast majority of popular LLMs are based on the Transformer structure. We selected the RWKV-7B model proposed in (Peng et al., 2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent. We analyzed this model using the methodology from 2.1. Figure 7 shows the redundancy of the RWKV-7B 111 We use rwkv-v5-world-7B from https://huggingface.co/BlinkDL/rwkv-5-world model. Through this figure, we can observe that the RWKV-7B model also exhibits a high level of redundancy. This may suggest that the redundancy is universal across LLMs.
@@ 4.4 Importance metric
Figure 8: Different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer
@@
Reverse-order/Norm: This metric posits that importance is inversely proportional to the sequence order. It assigns higher importance scores to the shallower layers Inspired by Figure 3. This method gives the same order as measuring importance by hidden states norm mentioned in LABEL:layerimportacne. Here, the importance metric is the layer’s index itself.
Relative Magnitude: Proposed in (Samragh et al., 2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
@@
Figure 8 demonstrates the different metrics. We observe that shallower layers in the LLM network are more crucial than deeper ones. Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics. The method of (Samragh et al., 2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent. It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation. This also indicates the limitations of the current LLMs evaluation methods.
@@ 5 Limitation
Although our method demonstrates strong competitiveness compared to current pruning methods, with some benchmarks even suggesting that deleting certain layers has no impact on the final result, this does not imply that layer removal comes without drawbacks. Due to the limitations of benchmarking, current evaluations may not fully capture the impact of layer removal on model performance. Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks. On benchmarks such as GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero. Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal. Post-training techniques could potentially further mitigate performance losses, which is an area of ongoing research.